# Agnostics: Learning to Code in Any Programming Language via Reinforcement with a Universal Learning Environment

### v1

@@ 6 Acknowledgments
This work was supported by Czech Science Foundation Grant No. 23-07580X. This work is partially supported by the U.S. National Science Foundation (SES-2326174). This material is based upon work supported by the U.S. Department of Energy, Office of Science under Award Number DESC0025613.
We used GNU parallel (Tange 2023).
@@
Appendix A Efficiency of Rejection Sampling
An alternative to reinforcement learning is to use rejection sampling with supervised fine-tuning: prompt the model to synthesize $n$ solutions to each task, reject solutions that fail tests, and fine-tune on the task-solution pairs that pass tests. If the model is weak at a task, which is the case with low-resource languages, then it is possible for all solutions to fail for a given task. Cassano et al. (2024) use this approach to get solutions for $\approx 30\%$ of the tasks in their dataset. They use $n=50$ for Lua and Julia, but increase to $n=100$ for OCaml.
The efficiency of this approach depends on the hardness of the task and the capabilities of the model. In this paper, we work with newer models that are marginally better at low-resource languages (based on MultiPL-E benchmark results). The task in Cassano et al. (2024) is to translate a simple, self-contained Python function from the model’s pretraining data into an equivalent function in another programming language. This task is significantly easier than the Agnostics task, which is to solve a competitive programming problem in a low-resource language, without any reference code or tests.
We can estimate the efficiency of rejection sampling from the rewards assigned in the early steps of reinforcement learning, when the policy is closest to the initial policy. Figure 7 shows the pass1 rates while training Qwen 3 4B on Fortran and OCaml. In the first 200 steps, the average pass rate for Fortran and OCaml are $1.9\%$ and $1.1\%$ respectively–considerably less than the $30\%$ success rate reported by Cassano et al. (2024) on their code translation task, making the rejection sampling approach prohibitively expensive for the low-resource language setting.
Appendix B Preparing Datasets for Agnostics
@@
is a dataset of Mostly Basic Programming Problems, transformed from the MBPP (Austin et al. 2021) dataset. The source dataset contains problems asking to complete a Python function from on its signature and a docstring, where the docstring specifies what result the function should return; our dataset contains equivalent programs which read/write the same data from/to standard input/output. Out of 974 problems, we are able to translate 776 problems into Ag-MBPP-X (348 problems are in the sanitized subset of MBPP).
@@
is a dataset of competitive programming problems, created from Codeforces problems in the open-r1/codeforces dataset. The source problems already specified programs by their I/O behavior, hence only very minor changes were needed to build language-universal Agnostics problems out of the fields in the dataset: we only skipped the time and memory restrictions present in the original problems. To be precise, we used data from the open-r1/codeforces-cots dataset, solutions_py_decontaminated subset, which contains problems decontaminated using 8-gram overlap against multiple benchmarks, in particular LiveCodeBench. We used the auxilliary checker_interactor subset to only keep the problems which admit a simple verifier for their solutions, i.e., problems where a single output is correct for each input and where the solution only needs to read data from its input, compute the result, and write it to the standard output. We prepared both a train and a test split. The former contains 5369 problems and the latter contains 107 problems we held out from the source dataset, 7 selected manually and 100 randomly. The train split features randomized prompts, which we found help with generalizing the results of training on the dataset to other benchmarks. The prompts were randomly split into a number of types. 30% of the prompts use standard Markdown headings to start different sections of the prompt, 35% use bold text instead, and the remaining 35% simply concatenate the prompt sections together. Half of the prompts of the final type also do not include an I/O sample in the prompt.
@@
is also a dataset of competitive programming problems, created from a subset of the LiveCodeBench dataset (Jain et al. 2024b). LiveCodeBench 5.0 has 880 problems, of which 381 have Python starter code and test cases. The remaining 499 problems do not use starter code and instead use standard I/O to specify and test solutions. Hence we used these problems to transform LiveCodeBench into an Agnostics dataset. Ag-LiveCodeBench-X only has a train split, like its source dataset.
Appendix C Agnostics Configurations
@@
The R configuration (the same as figure 4, repeated here for convenience):
@@
Use R version 4.
Use readLines(con = file("stdin"))‘ to
read input from stdin. Optionally, use
the n argument to read the first n
lines. For example:
@@
print(n) # print the first line of input
Also, use cat to print output to stdout.
For example:
@@
Always begin each scope with implicit none, pick explicit kinds via
selected_*_kind, and declare proper lengths-character(len=*) is legal only for
dummy arguments, not locals. Strings are blank-padded: call len_trim before
iterating, and store dynamic text in deferred-length allocatables
(character(len=:), allocatable :: s). List-directed read(*,*) arr does not
auto-size arrays; read a count first, then allocate and read, or tokenize a
line manually. When translating 0-based formulas (heaps, bit positions)
remember Fortran arrays default to 1-based; if you want 0-based, declare lower
bounds. Use real literals (2.0d0, 1.0_rk) to avoid silent integer division,
and guard against overflow when exponentiating integers. For frequency tables,
allocate an array or use findloc; Fortran lacks native dicts/sets, so you must
implement search yourself. Prefer array intrinsics (sum, count, pack) over
hand-rolled loops, and keep helper procedures inside a contains section or
module so interfaces are explicit. return inside the main program is
non-idiomatic; use structured blocks or stop. Never print interactive prompts
in batch solutions; just read, compute, and write.
@@
Appendix D Training and Results
D.1 Choosing hyperparameters
Before picking the hyperparameters described in § 3.3, we investigated other values by training the Qwen 3 4B model on a previous version of Ag-Codeforces-X. We trained two models for each of Lua, Julia and R, using a linear learning rate schedule with the same learning rate. We decided against it since some of the runs degraded the model’s capabilities, unlike any of the runs we did with a cosine decay schedule. We also compared between GRPO group sizes of 16, 32 and 64 by training the same model on Lua. We observed that in some runs with group size 16, the model improved significantly less than at higher group sizes. Since we found no significant difference between group sizes 32 and 64, we chose the smaller one due to limited resources.
D.2 Training dynamics
@@
In this section we discuss the measurements we took while training the models. Figure 8 shows the GRPO group batch pass@1 while training the Qwen3-4B-CF-X models. The scores of all the models are broadly correlated with one another, in part due to training them on the same permutation of the training data. Figures 10, 11, 12, 13 and 14 compare the GRPO group pass@1 scores with the pass@1 scores on the test split. We see that the scores on the test split is broadly correlated with the train split rewards. In most cases, we see that the train scores keep increasing until the end of the epoch, together with the test split pass@1 scores, indicating that the model keeps improving until the end of the dataset.
D.3 Hardware and Software Used
@@
Appendix E Bug Taxonomy
E.1 Prompt for Generating Taxonomy
@@
The prompt produced a taxonomy of 11 bug categories. We edited these categories and selected 7 categories relevant to us, shown in §E.2.
E.2 Bug Taxonomy Used For Analysis
@@
E.3 Radar Charts for All Programming Languages
@@
Appendix F Reproducibility Checklist
This paper:
  • Includes a conceptual outline and/or pseudocode description of AI methods introduced: yes
  • Clearly delineates statements that are opinions, hypothesis, and speculation from objective facts and results: yes
  • Provides well marked pedagogical references for less-familiar readers to gain background necessary to replicate the paper: yes
Does this paper make theoretical contributions? No.
Does this paper rely on one or more datasets? Yes.
  • A motivation is given for why the experiments are conducted on the selected datasets: yes.
  • All novel datasets introduced in this paper are included in a data appendix: no.
  • All novel datasets introduced in this paper will be made publicly available upon publication of the paper with a license that allows free usage for research purposes: yes.
  • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations: yes.
  • All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available: yes.
  • All datasets that are not publicly available are described in detail, with explanation why publicly available alternatives are not scientifically satisfying: N/A.
Does this paper include computational experiments? Yes.
  • This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting: yes.
  • Any code required for pre-processing data is included in the appendix: no.
  • All source code required for conducting and analyzing the experiments is included in a code appendix: no.
  • All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes: yes.
  • All source code implementing new methods have comments detailing the implementation, with references to the paper where each step comes from: no.
  • If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results: yes.
  • This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks: yes.
  • This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics: yes.
  • This paper states the number of algorithm runs used to compute each reported result: yes.
  • Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information: no.
  • The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank): no.
  • This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments: yes.

### v2

@@ Acknowledgments
This work is partially supported by the U.S. National Science Foundation (SES-2326174). This material is based upon work supported by the U.S. Department of Energy, Office of Science under Award Number DESC0025613.
This research used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy User Facility using NERSC award ALCC-ERCAP 0038272 (m5083-2024).
This work is supported by Czech Science Foundation Grant No. 23-07580X.
We thank Northeastern Research Computing for support with the Northeastern University Explorer cluster. This work used the Delta cluster at the National Center for Supercomputing Applications (NCSA) through allocation CIS230213 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants 2138259, 2138286, 2138307, 2137603, and 2138296.
This work uses vLLM (Kwon et al., 2023), Transformers (Wolf et al., 2020), and GNU Parallel (Tange, 2023).
@@
Appendix A Preparing Datasets for Agnostics
@@
is a dataset of Mostly Basic Programming Problems, transformed from the MBPP (Austin et al., 2021) dataset. The source dataset contains problems asking to complete a Python function from on its signature and a docstring, where the docstring specifies what result the function should return; our dataset contains equivalent programs which read/write the same data from/to standard input/output. Out of 974 problems, we are able to translate 776 problems into Ag-MBPP-X (348 problems are in the sanitized subset of MBPP). We analyzed Ag-MBPP-X for data similarity against Ag-LiveCodeBench-X using Decon (2025) and found no data overlap. The analysis was done based on 5-grams with no token sampling (see Decon documentation).
@@
is a dataset of competitive programming problems, created from Codeforces problems in the open-r1/codeforces dataset. The source problems already specified programs by their I/O behavior, hence only very minor changes were needed to build language-universal Agnostics problems out of the fields in the dataset: we only skipped the time and memory restrictions present in the original problems. To be precise, we used data from the open-r1/codeforces-cots dataset, solutions_py_decontaminated subset, which contains problems decontaminated using 8-gram overlap against multiple benchmarks, in particular LiveCodeBench. We used the auxilliary checker_interactor subset to only keep the problems which admit a simple verifier for their solutions, i.e., problems where a single output is correct for each input and where the solution only needs to read data from its input, compute the result, and write it to the standard output. We prepared both a train and a test split. The former contains 5369 problems and the latter contains 105 problems we held out from the source dataset, 5 selected manually and 100 randomly. The manually-selected problems were chosen to be much easier than average, to make it easier to detect if a model can solve any problems at all in a given programming language. In short, the 5 problems are: “output a number in binary notation”, “remove all digits from a string”, “check if the parentheses are balanced”, “parse two integers and add them”, and the following longer problem: “Petr stands in line of n people, but he doesn’t know exactly which position he occupies. He can say that there are no less than a people standing in front of him and no more than b people standing behind him. Find the number of different positions Petr can occupy.”
The train split features randomized prompts, which we found help with generalizing the results of training on the dataset to other benchmarks. The prompts were randomly split into a number of types. 30% of the prompts use standard Markdown headings to start different sections of the prompt, 35% use bold text instead, and the remaining 35% simply concatenate the prompt sections together. Most of the prompts follow the source dataset and feature an I/O sample in the prompt, with other samples withheld as private. Half of the prompts of the final type do not feature any I/O sample.
@@
is also a dataset of competitive programming problems, created from a subset of the LiveCodeBench dataset (Jain et al., 2024a). LiveCodeBench 5.0 has 880 problems, of which 381 have Python starter code and test cases. The remaining 499 problems do not use starter code and instead use standard I/O to specify and test solutions. Hence we used these problems to transform LiveCodeBench into an Agnostics dataset. Ag-LiveCodeBench-X only has a test split, like its source dataset.
Appendix B Agnostics Configurations
@@
The R configuration (unmodified, unlike figure 4):
@@
Use R version 4. Use readLines(con = file("stdin"))‘ to read input from stdin. Optionally, use the n argument to read the first n lines. For example:
@@
cat(n) # print the first line of input
Also, use cat to print output to stdout. For example:
@@
Always begin each scope with implicit none, pick explicit kinds via selected_*_kind, and declare proper lengths-character(len=*) is legal only for dummy arguments, not locals. Strings are blank-padded: call len_trim before iterating, and store dynamic text in deferred-length allocatables (character(len=:), allocatable :: s). List-directed read(*,*) arr does not auto-size arrays; read a count first, then allocate and read, or tokenize a line manually. When translating 0-based formulas (heaps, bit positions) remember Fortran arrays default to 1-based; if you want 0-based, declare lower bounds. Use real literals (2.0d0, 1.0_rk) to avoid silent integer division, and guard against overflow when exponentiating integers. For frequency tables, allocate an array or use findloc; Fortran lacks native dicts/sets, so you must implement search yourself. Prefer array intrinsics (sum, count, pack) over hand-rolled loops, and keep helper procedures inside a contains section or module so interfaces are explicit. return inside the main program is non-idiomatic; use structured blocks or stop. Never print interactive prompts in batch solutions; just read, compute, and write.
@@
Appendix C Training and Results
C.1 Choosing hyperparameters
Before picking the hyperparameters described in § 3.3, we investigated other values by training the Qwen 3 4B model on a previous version of Ag-Codeforces-X. We trained two models for each of Lua, Julia and R, using a linear learning rate schedule with the same learning rate. We decided against it since some of the runs degraded the model’s capabilities, unlike any of the runs we did with a cosine decay schedule.
We compared between GRPO group sizes of 16, 32 and 64 by training the same model on Lua. In some runs with group size 16, we saw the model improved significantly less than at higher group sizes. We ran an experiment to compared different temperature and group size settings (table 6).
Table 6: Hyperparameter sweep, pass@1 score.
Model Group size Temperature Ag-LiveCodeBench-X
Qwen3-4B-CF-Lua 32 0.7 23.00
normal-r1 32 0.7 19.87
normal-r2 32 0.7 21.58
size16-r1 16 0.7 19.80
size16-r2 16 0.7 19.40
size16-r3 16 0.7 19.65
size16-r4 16 0.7 20.61
size64-r1 64 0.7 21.71
size64-r2 64 0.7 21.21
size64-r3 64 0.7 20.87
temp0p2-r1 32 0.2 19.91
temp0p2-r2 32 0.2 20.39
temp0p2-r3 32 0.2 21.98
temp1-r1 32 1.0 21.22
temp1-r2 32 1.0 21.48
temp1-r3 32 1.0 20.30
The models trained at group size 16 had slightly lower scores compared to other models, while the ones trained at group size 64 displayed scores comparable to other models. However, they took significantly longer to train. Two group size 64 models took $\sim 20.5$h to train on average (the third one was trained on a different machine). In comparison, the group size 32 models trained at the same time on the same machine took $\sim 12$h on average. The models trained at temperatures other than the $0.7$ recommended by the Qwen team performed similarly to the other models.
As we found no significant difference between the temperature settings and between group sizes 32 and 64, we chose the smaller group size due to limited resources, and used the recommended temperature settings.
C.2 Training dynamics
@@
In this section we discuss the measurements we took while training the models. Figure 8 shows the GRPO group batch pass@1 while training the Qwen3-4B-CF-X models. The scores of all the models are broadly correlated with one another, which may at least in part be due to training them on the same permutation of the training data. Figures 10, 11, 12, 13 and 14 compare the GRPO group pass@1 scores with the pass@1 scores on the test split. We see that the scores on the test split is broadly correlated with the train split rewards. In most cases, we see that the train scores keep increasing until the end of the epoch, together with the test split pass@1 scores, indicating that the model keeps improving until the end of the dataset.
C.3 Reward function
We investigated the results of partial rewards. We trained Qwen 3 4B—the base model of Qwen3-4B-CF-X—on Ag-Codeforces-Lua, giving it a partial reward of $0.2$ if it generated code which failed one of the tests by producing wrong output but otherwise terminated without an error. The full reward for a snippet passing all tests was still 1. Table 7 shows that the trained models score below Qwen3-4B-CF-Fortran both on Ag-LiveCodeBench-Lua and on the test split of Ag-Codeforces-Lua (counting only the full-credit reward). The latter scores are particularly far lower, clearly showing that the models learned to abuse the partial-credit reward.
During training, we saw the models focus on the partial reward. The average result from the partial reward component was clearly increasing more quickly than the result from the full reward component. In the training generations we inspected, the models also often claimed to generate a “draft” answer and produced a program which ignored the problem in the prompt, for instance by only printing a hard-coded string such as “0”.
Table 7: Partial reward experiment, pass@1 rates.
Model Ag-LiveCodeBench-X Ag-Codeforces-X (test split)
Qwen3-4B-CF-Lua 23.00 24.76
partial-r1 18.57 13.33
partial-r2 20.16 15.62
C.4 Cross-Programming-Language Negative Transfer
To demonstrate that Agnostics training does not lower performance on different programming languages, we evaluated the models we trained on variants of Ag-LiveCodeBench-X (table 8).
Table 8: Cross-PL evaluation, pass@1 rates.
Model Ag-LiveCodeBench-X
X= Python Lua Julia R
Qwen 3 4B 34.34 11.00 10.00 10.00
[…]
Qwen 3 8B 33.51 11.00 9.00 9.00
[…]
DSC 6.7B Ins 16.86 7.89 4.79 7.97
[…]
C.5 Hardware and Software Used
@@
Appendix D Bug Taxonomy
D.1 Prompt for Generating Taxonomy
@@
The prompt produced a taxonomy of 11 bug categories. We edited these categories and selected 7 categories relevant to us, shown in §D.2.
D.2 Bug Taxonomy Used For Analysis
@@
D.3 Radar Charts for All Programming Languages